OpenStudy (anonymous):

How do I find the roots of 3x^2-7x+5

OpenStudy (jhannybean):

Have you learned the completing the square method?

OpenStudy (dtan5457):

Yep. Looks like your either gonna complete the square or use the quadratic formula, your polynomial can't evenly be factored

OpenStudy (mathstudent55):

\(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \)

OpenStudy (jhannybean):

@impossibledreams ?

OpenStudy (anonymous):

yes I have @Jhannybean

OpenStudy (jhannybean):

okay :) So first group your x terms together.\[(3x^2-7x)+5\]Now factor out a 3.\[3\left(x^2-\frac{7}{3}x\right)+5\]Now complete the quadratic inside the parenthesis. \[3\left(x^2-\frac{7}{3}x+\frac{49}{36}\right)+5-\frac{49}{12}\]Simplify. \[3\left(x-\frac{7}{6}\right)^2 -\frac{11}{12}\]To find the roots,set it equal to 0.\[3\left(x-\frac{7}{6}\right)^2 = \frac{11}{12}\]Isolate your function by dividing both sides by 3.\[\left(x-\frac{7}{6}\right)^2 =\frac{11}{36}\]square root both sides. \[x-\frac{7}{6} = \pm \frac{\sqrt{11}}{6}\]Solve for x.\[x=\frac{7 \pm \sqrt{11}}{6}\]

OpenStudy (dtan5457):

Oh my^

OpenStudy (dtan5457):

Quadratic formula looks easier in this case, lol.

OpenStudy (jhannybean):

I showed every step of the process, some of these steps are mental math where the function is easily simplified :)